# Comparison results for Markov tree distributions

### v1

@@ Appendix A Examples
The following example shows that the SI assumptions on $X$ and $Y$ in opposite order in Proposition 1.1 cannot be replaced by SI in the same order, i.e., there is no version of Proposition 1.1 under the assumptions that $X_{i+1}\uparrow_{st}X_{i}$ and $Y_{i+1}\uparrow_{st}Y_{i}\,,$ see also [45, Example 4.4]. We even show that, under two additional SI assumptions, see Figure 8 a) there is no supermodular comparison result.
Example A.3 (Condition (ii) in Proposition 1.1 cannot be replaced by $Y_{i+1}\uparrow_{st}Y_{i}$)
@@
$P[X_{i}=k,X_{i+1}=l]=a^{ii+1}_{kl}\text{ and }\quad P[Y_{i}=k,Y_{i+1}=l]=b^{ii+1}_{kl}\text{ for }k,l\in\{0,1,2\}\text{ and }i\in\{0,1\}\,.$
All marginal distributions of the vectors $X$ and $Y$ are uniform on $\{0,1,2\}\,,$ i.e., $X_{i}\stackrel{{\scriptstyle\mathrm{d}}}{{=}}Y_{i}\sim U(\{0,1,2\})$ for $i\in\{0,1,2\}\,.$ Further, for $i\in\{0,1\}\,,$ it holds that $(X_{i},X_{i+1})\leq_{lo}(Y_{i},Y_{i+1})$ and thus, due to identical marginals, $(X_{i},X_{i+1})\leq_{sm}(Y_{i},Y_{i+1})\,,$ see (9). Moreover,
@@
i.e., $X$ and $Y$ are CIS, see Lemma B.3, and the bivariate subvectors $(X_{1},X_{2})$ and $(Y_{1},Y_{2})$ are CI. However, $Y$ violates condition (ii) of Proposition 1.1 as well as condition (ii) of Theorem 1.3 since $Y_{0}\not\uparrow_{st}Y_{1}\,.$ For the lower orthant set $\Theta\coloneqq(-\infty,1]^{3}$, we obtain
@@
Example A.4 (Conditions (i)-(ii) in Lemma 1.2 cannot be replaced by $(X_{0},X_{i})$ CI and $Y_{0}\uparrow_{st}Y_{i}$)
@@
$\begin{split}a^{01}&=\frac{1}{40}\begin{pmatrix}3&3&3&1\\ 3&3&3&1\\ 3&3&3&1\\ 1&1&1&7\\ \end{pmatrix},\quad b^{01}=\frac{1}{40}\begin{pmatrix}4&3&2&1\\ 5&4&1&0\\ 1&2&5&2\\ 0&1&2&7\\ \end{pmatrix},\\ a^{02}&=\frac{1}{40}\begin{pmatrix}4&3&2&1\\ 3&3&2&2\\ 2&3&3&2\\ 1&1&3&5\\ \end{pmatrix},\quad b^{02}=\frac{1}{40}\begin{pmatrix}6&2&2&0\\ 3&3&1&3\\ 1&3&4&2\\ 0&2&3&5\\ \end{pmatrix}.\end{split}$ (41)
@@
$P[X_{0}=k,X_{i}=l]=a^{0i}_{kl}\text{ and }P[Y_{0}=k,Y_{i}=l]=b^{0i}_{kl}\text{ for }k,l\in\{0,1,2,3\}\text{ and }i\in\{0,1\}\,.$
All marginal distributions of the vectors $X$ and $Y$ are uniform on $\{0,1,2,3\}\,,$ i.e., $X_{i}\stackrel{{\scriptstyle\mathrm{d}}}{{=}}Y_{i}\sim U(\{0,1,2,3\})$ for $i\in\{0,1,2\}\,.$ Further, for $i\in\{1,2\}\,,$ it holds that $(X_{0},X_{i})\leq_{lo}(Y_{0},Y_{i})$ and thus, due to identical marginals, $(X_{0},X_{i})\leq_{sm}(Y_{0},Y_{i})\,,$ see (9). Moreover,
@@
The SI conditions in the following example correspond to the graph in Figure 8 c). This example is a comonotonic extension of Example A.4 and will also be used for the proof of Proposition 5.2 which provides the generality of the SI assumptions in Theorem 1.3.
Example A.5 (A commonotonic extension of Example A.4)
Let $(X_{0},X_{1},X_{2})$ and $(Y_{0},Y_{1},Y_{2})$ be the random vectors in Example A.4. Define the ($d+1$)-dimensional random vectors $X^{\prime}$ and $Y^{\prime}$ by
$X^{\prime}=(X^{\prime}_{0},\dots,X^{\prime}_{d})\coloneqq(X_{1},X_{0},\dots,X_{0},X_{2})\quad\text{and}\quad Y^{\prime}=(Y^{\prime}_{0},\dots,Y^{\prime}_{d})\coloneqq(Y_{1},Y_{0},\dots,Y_{0},Y_{2}).$
@@
$X^{\prime}_{i+1}$ $\uparrow_{st}X^{\prime}_{i},$ $X^{\prime}_{i}$ $\uparrow_{st}X^{\prime}_{i+1},$ for all $\text{for all }i\in\{0,\dots,d-1\},$
$Y^{\prime}_{i+1}$ $\uparrow_{st}Y^{\prime}_{i},$ $Y^{\prime}_{j}$ $\uparrow_{st}Y^{\prime}_{j+1},$ for all $\text{for all }i\in\{0,\dots,d-2\}\text{ and }j\in\{1,\dots,d-1\}.$
Moreover, it holds that $(X^{\prime}_{i},X^{\prime}_{i+1})\leq_{lo}(Y^{\prime}_{i},Y^{\prime}_{i+1})$ for all $i=0,\dots,d-1$. Using (42) we obtain for the lower orthant set $\Theta\coloneqq(-\infty,2]^{d+1}$ that $P^{X^{\prime}}(\Theta)>P^{Y^{\prime}}(\Theta).$ This shows that $X^{\prime}\nleq_{lo}Y^{\prime}$ and thus $X^{\prime}\nleq_{sm}Y^{\prime}$. This example highlights the importance of the assumption $Y_{j}\uparrow_{st}Y_{i}$ for all $(i,j)\in E$ with $j\notin P$ in Theorem 1.3 by extending the usage of Example A.4 to more complex cases, see also Proposition 5.2.
The following example shows that Proposition 5.4 based on the Schur order for conditional distributions cannot be extended to Markov processes and neither to Markov tree distributions.
Example A.6 (Schur order is not applicable to Markov processes)
@@
for all $k,\ell\in\{0,1,2\}\,,$ $i\in\{0,1,2,3\}$ and $s,t\in[0,1]\,.$ All marginal distributions of the vectors $X$ and $Y$ are uniform on $[0,3)\,,$ i.e., $X_{i}\stackrel{{\scriptstyle\mathrm{d}}}{{=}}Y_{i}\sim U([0,3))$ for all $i\in\{0,1,2,3\}\,.$ It holds that $(Y_{i},Y_{i+1})$ is $\mathrm{\text{MTP}_{2}}$ (thus, in particular, $Y_{i}\uparrow_{st}Y_{i+1}$ and $Y_{i+1}\uparrow_{st}Y_{i}$, see (14)) and $(X_{i},X_{i+1})\leq_{lo}(Y_{i},Y_{i+1})$ for all $i\in\{0,1,2\}$. Further, the bivariate distributions fulfill the Schur order criteria
@@
$P^{X}(\Theta)=9\!\!\!\sum_{i_{1},\dots,i_{4}=1}^{2}a_{i_{1}i_{2}}a_{i_{2}i_{3}}a_{i_{3}i_{4}}=\dfrac{1.259}{3}>\dfrac{1.256}{3}=9\!\!\!\sum_{i_{1},\dots,i_{4}=1}^{2}b_{i_{1}i_{2}}b_{i_{2}i_{3}}b_{i_{3}i_{4}}=P^{Y}(\Theta),$
@@
which shows that $X\nleq_{lo}Y$ and consequently $X\nleq_{sm}Y$. Note that Theorem 1.3 cannot be applied because $X_{i}\not\uparrow_{st}X_{j}$ so that the vector $X$ violates assumption (i).
Hence, the random vector $X=(X_{0},\ldots,X_{3})$ is not smaller in supermodular order than $Y=(Y_{0},\ldots,Y_{3})$ even though, for all edges $(i,j)\in E\,,$ on the one hand $(X_{i},X_{j})$ has less variability and is closer to independence than $(Y_{i},Y_{j})$ with respect to the Schur order for conditional distributions, see (44). On the other hand, $(X_{i},X_{j})$ is PSMD and $(Y_{i},Y_{j})$ is $\mathrm{\text{MTP}_{2}}\,,$ i.e., all bivariate specifications exhibit positive dependencies.
@@
To summarize, this example shows the importance of the SI assumptions in Theorem 1.3 on both $(X_{n})_{n\in N}$ and $(Y_{n})_{n\in N}$ noting that the SI assumptions on $(X_{n})_{n\in N}$ cannot be skipped or weakened to PSMD even under the stronger comparison of the bivariate dependence specifications with respect to the Schur order for conditional distributions. This example also emphasises the special properties of star structures which allow a more general supermodular comparison result based on the Schur order for conditional distributions, see Proposition 5.4.
Appendix B Proofs of Section 1

### v3

@@ 6 Counterexamples
The following example shows that the SI assumptions on $X$ and $Y$ in ’opposite order’ in Proposition 1.1 cannot be replaced by SI in the ’same order’, i.e., there is no version of Proposition 1.1 under the assumptions that $X_{i+1}\uparrow_{st}X_{i}$ and $Y_{i+1}\uparrow_{st}Y_{i}\,;$ see also [33, Example 4.4]. More specifically, we show that even with two additional SI assumptions, as illustrated in Figure 8 a), no supermodular comparison result can be established.
Example 6.1 (Condition (ii) in Proposition 1.1 cannot be replaced by $Y_{i+1}\uparrow_{st}Y_{i}$)
@@
$P[X_{i}=k,X_{i+1}=l]=a^{ii+1}_{kl}\text{ and }P[Y_{i}=k,Y_{i+1}=l]=b^{ii+1}_{kl}$
for $k,l\in\{0,1,2\}$ and $i\in\{0,1\}.$ All marginal distributions of the vectors $X$ and $Y$ are uniform on $\{0,1,2\},$ i.e., $X_{i}\stackrel{{\scriptstyle\mathrm{d}}}{{=}}Y_{i}\sim U(\{0,1,2\})$ for $i\in\{0,1,2\}.$ Further, for $i\in\{0,1\},$ it holds that $(X_{i},X_{i+1})\leq_{lo}(Y_{i},Y_{i+1})$ and thus, due to identical marginals, $(X_{i},X_{i+1})\leq_{sm}(Y_{i},Y_{i+1})\,;$ see (6). Moreover,
@@
i.e., $X$ and $Y$ are CIS (see Lemma B.3), and the bivariate subvectors $(X_{1},X_{2})$ and $(Y_{1},Y_{2})$ are CI. However, $Y$ violates condition (ii) of Proposition 1.1 as well as condition (ii) of Theorem 1.3 since $Y_{0}\not\uparrow_{st}Y_{1}.$ For the lower orthant set $\Theta\coloneqq(-\infty,1]^{3}$, we obtain
@@
[…]
Figure 8: The graphs illustrate the non-sufficient SI conditions in a) Example 6.1, b) Example 6.2, c) Example 6.3 and d) Example 6.6, where an arrow $U\to V$ indicates $V\uparrow_{st}U$ and where $U\xleftrightarrow{\mathrm{\text{MTP}_{2}}}V$ indicates that $(U,V)$ is $\mathrm{\text{MTP}_{2}}$ which implies, in particular, that $U\uparrow_{st}V$ and $V\uparrow_{st}U.$ As shown in the respective examples, none of the SI conditions on $X=(X_{n})_{n\in N}$ and $Y=(Y_{n})_{n\in N}$ are sufficient for $X\leq_{sm}Y$ whenever $(X_{i},X_{j})\leq_{sm}(Y_{i},Y_{j})$ for all $(i,j)\in E.$
@@
Example 6.2 (Conditions (i)-(ii) in Lemma 1.2 cannot be replaced by $(X_{0},X_{i})$ CI and $Y_{0}\uparrow_{st}Y_{i}$)
@@
[…]
@@
$P[X_{0}=k,X_{i}=l]=a^{0i}_{kl}\text{ and }P[Y_{0}=k,Y_{i}=l]=b^{0i}_{kl}$
for $k,l\in\{0,1,2,3\}$ and $i\in\{0,1\}.$ All marginal distributions of the vectors $X$ and $Y$ are uniform on $\{0,1,2,3\},$ i.e., $X_{i}\stackrel{{\scriptstyle\mathrm{d}}}{{=}}Y_{i}\sim U(\{0,1,2,3\})$ for $i\in\{0,1,2\}.$ Further, for $i\in\{1,2\},$ it holds that $(X_{0},X_{i})\leq_{lo}(Y_{0},Y_{i})$ and thus, due to identical marginals, $(X_{0},X_{i})\leq_{sm}(Y_{0},Y_{i})\,;$ see (6). Moreover,
@@
The SI conditions in the following example correspond to the graph in Figure 8 c). This example is a comonotonic extension of Example 6.2 and is also used in the proof of Proposition A.2 on the generality of the SI assumptions in Theorem 1.3.
Example 6.3 (A comonotonic extension of Example 6.2)
Let $(X_{0},X_{1},X_{2})$ and $(Y_{0},Y_{1},Y_{2})$ be the random vectors in Example 6.2. Define the ($d+1$)-dimensional random vectors $X^{\prime}$ and $Y^{\prime}$ by
$X^{\prime}$ $=(X^{\prime}_{0},\dots,X^{\prime}_{d})\coloneqq(X_{1},X_{0},\dots,X_{0},X_{2})\quad\text{ and }$
$Y^{\prime}$ $=(Y^{\prime}_{0},\dots,Y^{\prime}_{d})\coloneqq(Y_{1},Y_{0},\dots,Y_{0},Y_{2}).$
@@
$X^{\prime}_{i+1}$ $\uparrow_{st}X^{\prime}_{i},$ $X^{\prime}_{i}$ $\uparrow_{st}X^{\prime}_{i+1},$ $\text{for all }i\in\{0,\dots,d-1\},$
$Y^{\prime}_{i+1}$ $\uparrow_{st}Y^{\prime}_{i},$ $Y^{\prime}_{j}$ $\uparrow_{st}Y^{\prime}_{j+1},$ $\text{for all }i\in\{0,\dots,d-2\}\text{ and }j\in\{1,\dots,d-1\}.$
Moreover, it holds that $(X^{\prime}_{i},X^{\prime}_{i+1})\leq_{lo}(Y^{\prime}_{i},Y^{\prime}_{i+1})$ for all $i=0,\dots,d-1$. Using (37) we obtain for the lower orthant set $\Theta\coloneqq(-\infty,2]^{d+1}$ that $P^{X^{\prime}}(\Theta)>P^{Y^{\prime}}(\Theta).$ This shows that $X^{\prime}\nleq_{lo}Y^{\prime}$ and thus $X^{\prime}\nleq_{sm}Y^{\prime}$. This example highlights the importance of the assumption $Y_{j}\uparrow_{st}Y_{i}$ for all $(i,j)\in E$ with $j\notin P$ in Theorem 1.3 by extending the usage of Example 6.2 to more complex cases; see also Proposition A.2.
The following example shows that Theorem 3.1 fails to hold if the assumption on equal ranges of the marginal distribution functions is not imposed. Additionally, the example highlights that the dependence structure of a random vector determined by a Markov realization of a bivariate tree specification is determined not only by the set of bivariate copulas and the conditional independence assumptions but also by the choice of the marginal distributions.
Example 6.4 (General marginals are not sufficient for Theorem 3.1)
$F_{0}=F_{1}=F_{2}=G_{1}=G_{2}=F_{U(0,1)}\quad\text{and}\quad G_{0}=\mathds{1}_{[0,\infty)}.$
Then $F_{0},F_{1},F_{2},G_{1}$ and $G_{2}$ are continuous distribution functions and $G_{0}$ is the distribution function of the Dirac distribution in $0,$ which is not continuous. Denote by $M$ the bivariate comonotonicity copula, i.e., $M(u,v)=\min\{u,v\}$ for $(u,v)\in[0,1],$ and consider the bivariate specifications $B_{e}:=C_{e}:=M$ for $e\in E.$ Let $U$ and $V$ be independent and uniformly on $(0,1)$ distributed random variables. Then, for $F=(F_{0},F_{1},F_{2}),G=(G_{0},G_{1},G_{2}),$ $B=(B_{e})_{e\in E},$ and $C=(C_{e})_{e\in E},$ the random vectors $X=(X_{0},X_{1},X_{2}):=(U,U,U)$ and $Y=(Y_{0},Y_{1},Y_{2}):=(0,U,V)$ are Markov realizations of $(F,T,B)$ and $(G,T,C),$ respectively, i.e.,
It holds that $F_{i}\geq_{st}G_{i}$ for all $i\in\{0,1,2\}$ and $B_{e}=C_{e}$ (which trivially implies $B_{e}\leq_{lo}C_{e}$) for $e\in E.$ Further, the bivariate copulas $B_{e}$ and $C_{e},$ $e\in E,$ are CI. Hence, all assumptions of Theorem 3.1(ii) except of continuity of the marginal distributions are satisfied. However, it holds for all $(u,v)\in(0,1)$ that
$F_{X_{1},X_{2}}(u,v)=\min\{u,v\}>uv=F_{Y_{1},Y_{2}}(u,v),$
which implies $(X_{1},X_{2})\nleq_{dsm}(Y_{1},Y_{2}).$ Since the decreasing supermodular order is closed with respect to marginalization, it follows that $X\nleq_{dsm}Y$.
Replacing in the above setting $G_{0}$ by $G_{0}=\mathds{1}_{[1,\infty)}$ and considering $Y=(1,U,V),$ it follows similarly that $(X_{1},X_{2})\nleq_{ism}(Y_{1},Y_{2})$ and thus $X\nleq_{ism}Y.$ Hence, we conclude that the continuity assumption for the marginal specifications in Theorem 3.1(i) cannot be omitted.
Note that in this example, $X_{1}$ and $X_{2}$ are comonotonic while $Y_{1}$ and $Y_{2}$ are independent. Since the bivariate tree specifications $(F,T,B)$ and $(G,T,C)$ only differ in the first component of the marginal specifications (i.e., $F_{0}\neq G_{0},$ $F_{1}=G_{1},$ $F_{2}=G_{2},$ and $B=C$), we conclude that the marginal distributions can also influence the dependence structure of Markov tree distributions. In other words, conditional independence is not solely a copula-based property.
The following example justifies the $\mathrm{\text{MTP}_{2}}$ assumption in Theorem 3.5 since, surprisingly, bivariate CI specifications generally do not result in a Markov realization that is CI.
Example 6.5 (A non-CI Markov tree distribution with bivariate CI marginals)
$P[X_{i}=k]$ $=1/3\quad\text{for }i,k\in\{0,1,2\},$ (38)
$P[X_{i}=k,X_{i+1}=\ell]$ $=a^{i+1}_{k\ell}\quad\text{for }i\in\{0,1\}\text{ and }k,\ell\in\{0,1,2\}.$ (39)
Note that the univariate distributions in (38) are the marginals of the bivariate distributions in (39). From the definition of the matrices $a^{1}$ and $a^{2}$ it can be seen that the subvectors $(X_{0},X_{1})$ and $(X_{1},X_{2})$ are CI . It even follows that $(X_{0},X_{2})$ is CI because, as a consequence of Lemma B.3, $X_{0}\uparrow_{st}(X_{1},X_{2})$ and $X_{2}\uparrow_{st}(X_{1},X_{0}).$ However, the vector $(X_{0},X_{1},X_{2})$ is not CI because
The following example shows that Proposition 4.2 based on the Schur order for conditional distributions cannot be extended to Markov processes and thus neither to Markov tree distributions.
Example 6.6 (Schur order is not applicable to Markov processes)
@@
for all $k,\ell\in\{0,1,2\},$ $i\in\{0,1,2,3\}$ and $s,t\in[0,1].$ All marginal distributions of the vectors $X$ and $Y$ are uniform on $[0,3),$ i.e., $X_{i}\stackrel{{\scriptstyle\mathrm{d}}}{{=}}Y_{i}\sim U([0,3))$ for all $i\in\{0,1,2,3\}.$ It holds that $(Y_{i},Y_{i+1})$ is $\mathrm{\text{MTP}_{2}}$ (thus, in particular, $Y_{i}\uparrow_{st}Y_{i+1}$ and $Y_{i+1}\uparrow_{st}Y_{i}$; see (8)) and $(X_{i},X_{i+1})\leq_{lo}(Y_{i},Y_{i+1})$ for all $i\in\{0,1,2\}$. Further, the bivariate distributions satisfy the Schur order
@@
$P^{X}(\Theta)=9\!\!\!\sum_{i_{1},\dots,i_{4}=1}^{2}a_{i_{1}i_{2}}a_{i_{2}i_{3}}a_{i_{3}i_{4}}$ $=\dfrac{1.259}{3}$
$>\dfrac{1.256}{3}=9\!\!\!\sum_{i_{1},\dots,i_{4}=1}^{2}b_{i_{1}i_{2}}b_{i_{2}i_{3}}b_{i_{3}i_{4}}=P^{Y}(\Theta),$
@@
which shows that $X\nleq_{lo}Y$ and, consequently, $X\nleq_{sm}Y$. Note that Theorem 1.3 cannot be applied because $X_{i}\not\uparrow_{st}X_{j}$ so that the vector $X$ violates assumption (i).
@@
Hence, the random vector $X=(X_{0},\ldots,X_{3})$ is not smaller in supermodular order than $Y=(Y_{0},\ldots,Y_{3})$ even though, for all edges $(i,j)\in E,$ $(X_{i},X_{j})$ has less variability and is closer to independence than $(Y_{i},Y_{j})$ with respect to the Schur order for conditional distributions; see (41). Remarkably, $(X_{i},X_{j})$ is PSMD and $(Y_{i},Y_{j})$ is $\mathrm{\text{MTP}_{2}},$ so that all bivariate specifications exhibit positive dependencies. To summarize, this example highlights the critical role of the SI assumptions in Theorem 1.3 for both $(X_{n})_{n\in N}$ and $(Y_{n})_{n\in N}.$ In particular, it shows that the SI assumptions on $(X_{n})_{n\in N}$ cannot be omitted or weakened to PSMD, even under the stronger comparison of the bivariate dependence specifications with respect to the Schur order for conditional distributions. Additionally, this example emphases the special properties of star structures, which enable a more general supermodular comparison result based on the Schur order for conditional distributions, as stated in Proposition 4.2.
Appendix A Discussion of the SI assumptions in Theorem 1.3
Our main result, Theorem 1.3, provides simple conditions for a supermodular comparison of Markov tree distributions. At first glance, the SI conditions on $X$ and $Y$ may appear rather unintuitive. As we demonstrate in the sequel, under the necessary ordering assumption (iii) of Theorem 1.3, these conditions cannot be omitted or weakened to PSMD.
Figure 8 illustrates sufficient SI conditions in the three-dimensional case which, together with ordering assumption (iii), lead to the supermodular comparison of Markov tree distributions. Whenever an SI condition in Theorem 1.3 is skipped, it results in the existence of subvectors with a dependence structure that aligns with (or is even weaker than) the SI conditions depicted in the graphs in Figure 8 a),b), or c). For these settings, we have provided examples that do not yield a lower orthant comparison result and thus do not imply a supermodular comparison result either. Proposition A.2 summarizes the findings of this section regarding the necessity of the SI conditions. In conclusion, for a Markov tree distribution on $3$ nodes, it is precisely these two SI conditions in Figure 8 c) that lead to a supermodular comparison result.
The following remark summarizes Examples 6.16.3 demonstrating that the SI conditions in Figure 8 do not yield comparison results as stated in Proposition 1.1 and Lemma 1.2.
Remark A.1
  1. (a)
    Condition (ii) in Proposition 1.1 cannot be replaced by $Y_{i+1}\uparrow_{st}Y_{i}:$ As shown in Example 6.1, there exist Markov tree distributed random vectors $X=(X_{0},X_{1},X_{2})$ and $Y=(Y_{0},Y_{1},Y_{2})$ for $T$ satisfying the SI conditions in Figure 8 a) such that $(X_{i},X_{i+1})\leq_{sm}(Y_{i},Y_{i+1})$ but $X\not\leq_{lo}Y.$ Hence, the SI conditions in Figure 8 a) do not imply a supermodular comparison result.
  2. (b)
    Conditions (i)-(ii) in Lemma 1.2 cannot be replaced by $(X_{0},X_{i})$ being CI and $Y_{0}\uparrow_{st}Y_{i}:$ As shown in Example 6.2, there exist Markov tree distributed random vectors $X=(X_{0},X_{1},X_{2})$ and $Y=(Y_{0},Y_{1},Y_{2})$ for $T$ satisfying the SI conditions in Figure 8 b) such that $(X_{0},X_{i})\leq_{sm}(Y_{0},Y_{i})$ for $i\in\{1,2\}$ but $X\not\leq_{lo}Y.$
  3. (c)
    As an extension of the setting in Figure 8 b), the SI conditions in Figure 8 c) are not sufficient to prove a supermodular ordering result as in Proposition 1.1; see Example 6.3.
Using Examples 6.16.3 summarized in the preceding remark, we show in the following proposition that none of the SI conditions in assumptions (i) and (ii) of Theorem 1.3 can be dropped or weakened to PSMD. Since the supermodular order is closed under marginalization (see [69, Theorem 9.A.9(c)]), the proof reduces to the cases considered above.
To this end, let $T=(N,E)$ be the tree, $P$ the path, and $k^{*}$ the specific child of the root as considered in Theorem 1.3. For a node $j^{*}\in N\backslash\{0\}$, which will be specified in Proposition A.2, we relax the SI conditions in (i) and (ii) of Theorem 1.3 concerning only one edge: For $(i,j)\in E$ let
$X_{j}\uparrow_{st}X_{i}\text{ if }j\notin\{k^{*},j^{*}\}\text{, and }(X_{p_{j^{*}}}^{\perp},X_{j^{*}}^{\perp})\leq_{sm}(X_{p_{j^{*}}},X_{j^{*}}),$ ($i^{*}$)
$Y_{i}\uparrow_{st}Y_{j}\text{ if }j\notin L\cup\{j^{*}\}\text{, }Y_{j}\uparrow_{st}Y_{i}\text{ if }j\notin P\text{, and }(X_{p_{j^{*}}}^{\perp},X_{j^{*}}^{\perp})\leq_{sm}(X_{p_{j^{*}}},X_{j^{*}}),$ ($ii^{*}$)
$Y_{i}\uparrow_{st}Y_{j}\text{ if }j\notin L\text{, }Y_{j}\uparrow_{st}Y_{i}\text{ if }j\notin P\cup\{j^{*}\}\text{, and }(X_{p_{j^{*}}}^{\perp},X_{j^{*}}^{\perp})\leq_{sm}(X_{p_{j^{*}}},X_{j^{*}}).$ ($ii^{**}$)
Note that $p_{j^{*}}\in N$ denotes the parent node of $j^{*}\in N$; see Definition B.1. We distinguish between three cases. The first one concerns the SI conditions on $X$ in (i). The second case relates to the SI conditions (ii) on $Y$, which states that $Y_{i}\uparrow_{st}Y_{j}$ for $(i,j)\in E$ whenever $j\notin L\cup\{0\}$, where $N\backslash(L\cup\{0\})$ is a nonempty set if and only if $T$ is not a star. In the third case we consider the SI conditions of the second part of (ii), which asserts that $Y_{j}\uparrow_{st}Y_{i}$ for $(i,j)\in E$, whenever $j\notin P\cup\{0\}$, where $N\backslash(P\cup\{0\})$ is a nonempty set if and only if $T$ is not a chain. The following result establishes that none of the SI conditions in Theorem 1.3 can be omitted.
Proposition A.2 (Generality of the SI assumptions of Theorem 1.3)
Assume that $|N|\geq 3.$ None of the SI assumptions in Theorem 1.3 can be replaced by the weaker concept of PSMD, i.e., for the tree $T=(N,E)$, the following statements hold true:
  1. (a)
    For any $j^{*}\in N\setminus\{0,k^{*}\}$, there are Markov tree distributed sequences $X=(X_{n})_{n\in N}$ and $Y=(Y_{n})_{n\in N}$ with $X\nleq_{lo}Y$ such that $X$ and $Y$ fulfill for all $e=(i,j)\in E$ condition ($i^{*}$) and conditions (ii) and (iii) of Theorem 1.3.
  2. (b)
    If $T$ is not a star, then for any $j^{*}\in N\setminus(L\cup\{0\})$, there are Markov tree distributed sequences $X=(X_{n})_{n\in N}$ and $Y=(Y_{n})_{n\in N}$ with $X\nleq_{lo}Y$, such that $X$ and $Y$ fulfill for all $e=(i,j)\in E$ condition ($ii^{*}$) and conditions (i) and (iii) of Theorem 1.3.
  3. (c)
    If $T$ is not a chain, then for any $j^{*}\in N\setminus(P\cup\{0\})$, there are Markov tree distributed sequences $X=(X_{n})_{n\in N}$ and $Y=(Y_{n})_{n\in N}$ with $X\nleq_{lo}Y$, such that $X$ and $Y$ fulfill for all $e=(i,j)\in E$ condition ($ii^{**}$) and conditions (i) and (iii) of Theorem 1.3.
In particular, in each of the statements (a)-(c), it holds that $X\not\leq_{sm}Y$.
Appendix B Proofs
B.1 Proofs of Section 1